### v1

@@ 1.7 Radiative Transfer Modeling
| Telescope/Instrument/Filter | Date |
| Danish-1.54m/DFOSC/Za | 2013-06-03 | ||||||
| VLT/Hawk-I/Ksa | 2014-02-16 | ||||||
| SAAO/SHOCa | 2014-04-29 | ||||||
| Springbok/Raptor/Emptya | 2014-04-29 | ||||||
| Hakos/Raptor/Clearb | 2017-04-09 | - | |||||
| WabiLodge/Raptor/Clearb | 2017-04-09 | ||||||
| WeaverRL/Raptor/Clearb | 2017-04-09 | ||||||
| Tivoli/Raptor/Clearb | 2017-06-22 | ||||||
| WdhBackes/Raptor/Clearb | 2017-06-22 | ||||||
| Onduruquea/Raptor/Clearb | 2017-06-22 | ||||||
| Outeniqua/Raptor/Clearb | 2017-06-22 | ||||||
| WdhMeza/ASI178MM/Clearb | 2017-06-22 | ||||||
| Hakos/Raptor/Clearb | 2017-06-22 | ||||||
| Danish-1.54m/Redb | 2017-07-23 | ||||||
| VLT/Hawk-I/Ksb | 2017-07-23 | ||||||
| SSO/Andor/g/I+zb | 2017-07-23 | ||||||
| ElRodeo/MerlinKite/Clearb | 2017-07-23 | ||||||
| ElSalvador/Raptor/Clearb | 2017-07-23 | ||||||
| Danish-1.54m/Visualb | 2017-07-23 | ||||||
| LaSilla1m/Raptor/Clearb | 2017-07-23 | ||||||
| OPD/Andor/Clearb | 2017-07-23 | ||||||
| TolarGrande/Raptor/Clearb | 2017-07-23 | ||||||
| IncaDeOro/Raptor/Clearb | 2017-07-23 | ||||||
| OPD/Andor/Clearb | 2017-08-24 | ||||||
| Glenlee/Watec/Clearb | 2020-06-19 | ||||||
| JWST/NIRCam/F150W2c | 2022-10-18 | ||||||
| JWST/NIRCam/F322W2c | 2022-10-18 |
@@
| Telescope/Instrument/Filter | Date | |||||||
| (m) | (km) | (km) | (km) | (km) | ||||
| Danish-1.54m/DFOSC/Za | 03/06/2013 | 1.054 | ||||||
| Springbok/Raptor/Emptya | 29/04/2014 | 0.565 | ||||||
| Gifberg/Raptor/Emptya | 29/04/2014 | 0.565 | ||||||
| Danish-1.54m/Visualb | 23/07/2017 | 0.510 | ||||||
| Danish-1.54m/Redb | 23/07/2017 | 0.770 | ||||||
| OPD/Andor/Clearb | 23/07/2017 | 0.640 | ||||||
| VLT/Hawk-I/K | 23/07/2017 | 2.145 | ||||||
| JWST/NIRCam/F150W2c | 18/10/2022 | 1.655 | ||||||
| JWST/NIRCam/F322W2c | 18/10/2022 | 3.225 | - | - | - | - |
| Telescope/Instrument/Filter | Date | |||||||
| (m) | (km) | (km) | (km) | (km) | ||||
| Danish-1.54m/DFOSC/Za | 03/06/2013 | 1.054 | ||||||
| Springbok/Raptor/Emptya | 29/04/2014 | 0.565 | ||||||
| Gifberg/Raptor/Emptya | 29/04/2014 | 0.565 | ||||||
| Danish-1.54m/Visualb | 23/07/2017 | 0.510 | ||||||
| Danish-1.54m/Redb | 23/07/2017 | 0.770 | ||||||
| OPD/Andor/Clearb | 23/07/2017 | 0.640 | ||||||
| VLT/Hawk-I/Ksb | 23/07/2017 | 2.145 | ||||||
| JWST/NIRCam/F150W2c | 18/10/2022 | 1.655 | ||||||
| JWST/NIRCam/F322W2c | 18/10/2022 | 3.225 |
@@ Acknowledgements
This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program #1271. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular, the institutions participating in the Gaia Multilateral Agreement. P.S-S. acknowledges financial support from the Spanish I+D+i project PID2022-139555NB-I00 (TNO-JWST) funded by MCIN/AEI/10.13039/501100011033. Y.K. is forever thankful to TÜBİTAK National Observatory (1997-2024) for its unwavering support of his career and science. The authors acknowledge financial support from the Severo Ochoa grant CEX2021-001131-S funded by MICIU/AEI/ 10.13039/501100011033. M.A. thanks CNPq grants 427700/2018-3, 310683/2017-3 and 473002/2013-2. A.R.G.J. thanks the financial support of São Paulo Research Foundation (FAPESP, proc. 2016/24561-0 and 2018/11239-8). C.L.P. is thankful for the support of the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brazil (CAPES) and Fundação de Amparo à Pesquisa do Estado do Rio de Janeiro - FAPERJ/DSC-10 (E26/204.141/2022). HBH and SNM also acknowledge support from NASA JWST Interdisciplinary Scientist grant 21-SMDSS21-0013.

### v2

@@ 1.7 Radiative Transfer Modeling
References.: a. [Berard2017], b. [Morgado2021], c. This work.
| Telescope/Instrument/Filter | Date |
| Danish-1.54m/DFOSC/Za | 2013-06-03 | ||||||||||||
| VLT/Hawk-I/Ksa | 2014-02-16 | ||||||||||||
| SAAO/SHOCa | 2014-04-29 | ||||||||||||
| Springbok/Raptor/Emptya | 2014-04-29 | ||||||||||||
| Hakos/Raptor/Clearb | 2017-04-09 | – | – | ||||||||||
| WabiLodge/Raptor/Clearb | 2017-04-09 | ||||||||||||
| WeaverRL/Raptor/Clearb | 2017-04-09 | ||||||||||||
| Tivoli/Raptor/Clearb | 2017-06-22 | ||||||||||||
| WdhBackes/Raptor/Clearb | 2017-06-22 | ||||||||||||
| Onduruquea/Raptor/Clearb | 2017-06-22 | ||||||||||||
| Outeniqua/Raptor/Clearb | 2017-06-22 | ||||||||||||
| WdhMeza/ASI178MM/Clearb | 2017-06-22 | ||||||||||||
| Hakos/Raptor/Clearb | 2017-06-22 | ||||||||||||
| Danish-1.54m/Redb | 2017-07-23 | ||||||||||||
| VLT/Hawk-I/Ksb | 2017-07-23 | ||||||||||||
| SSO/Andor/g/I+zb | 2017-07-23 | ||||||||||||
| ElRodeo/MerlinKite/Clearb | 2017-07-23 | ||||||||||||
| ElSalvador/Raptor/Clearb | 2017-07-23 | ||||||||||||
| Danish-1.54m/Visualb | 2017-07-23 | ||||||||||||
| LaSilla1m/Raptor/Clearb | 2017-07-23 | ||||||||||||
| OPD/Andor/Clearb | 2017-07-23 | ||||||||||||
| TolarGrande/Raptor/Clearb | 2017-07-23 | ||||||||||||
| IncaDeOro/Raptor/Clearb | 2017-07-23 | ||||||||||||
| OPD/Andor/Clearb | 2017-08-24 | ||||||||||||
| Glenlee/Watec/Clearb | 2020-06-19 | ||||||||||||
| JWST/NIRCam/F150W2c | 2022-10-18 | ||||||||||||
| JWST/NIRCam/F322W2c | 2022-10-18 |
References.: a. [Berard2017], b. [Morgado2021], c. This work.
| Telescope/Instrument/Filter | Date |
| Danish-1.54m/DFOSC/Za | 03/06/2013 | ||||||||||||
| Springbok/Raptor/Emptya | 29/04/2014 | ||||||||||||
| Gifberg/Raptor/Emptya | 29/04/2014 | ||||||||||||
| Danish-1.54m/Visualb | 23/07/2017 | ||||||||||||
| Danish-1.54m/Redb | 23/07/2017 | ||||||||||||
| OPD/Andor/Clearb | 23/07/2017 | ||||||||||||
| VLT/Hawk-I/K | 23/07/2017 | ||||||||||||
| JWST/NIRCam/F150W2c | 18/10/2022 |
References.: a. [Berard2017], b. [Morgado2021], c. This work.
@@ Acknowledgements
This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program #1271. All the JWST data used in this paper can be found in MAST: https://doi.org/10.17909/qaaf-8q45. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular, the institutions participating in the Gaia Multilateral Agreement. P.S-S. acknowledges financial support from the Spanish I+D+i project PID2022-139555NB-I00 (TNO-JWST) funded by MCIN/AEI/10.13039/501100011033. Y.K. is forever thankful to TÜBİTAK National Observatory (1997-2024) for its unwavering support of his career and science. The authors acknowledge financial support from the Severo Ochoa grant CEX2021-001131-S funded by MICIU/AEI/ 10.13039/501100011033. M.A. thanks CNPq grants 427700/2018-3, 310683/2017-3 and 473002/2013-2. A.R.G.J. thanks the financial support of São Paulo Research Foundation (FAPESP, proc. 2016/24561-0 and 2018/11239-8). C.L.P. is thankful for the support of the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brazil (CAPES) and Fundação de Amparo à Pesquisa do Estado do Rio de Janeiro - FAPERJ/DSC-10 (E26/204.141/2022). HBH and SNM also acknowledge support from NASA JWST Interdisciplinary Scientist grant 21-SMDSS21-0013. CK was supported by the K-138962 and TKP2021-NKTA-64 grants of the National Research, Development and Innovation Office (NKFIH), Hungary. We sincerely thank both anonymous referees for their exceptional and insightful contributions, which significantly improved the clarity and quality of this manuscript.